In a family, a couple has a son and a daughter. The age of the father is four times that of his daughter and the age of the son is half of that of his mother. The wife is 8 years younger to her husband and the brother is seven years older than his sister. What is the age of the mother?
Answer by ptfile(81)

The age of the father is four times that of his daughter.
f=4s
the age of the son is half of that of his mother.
b=1/2m or m=2b
brother is seven years older than his sister.
b=s+7
The wife is 8 years younger to her husband.
f=m+8 or 4s=2(s+7)+8
4s=2(s+7)+8
4s=2s+14+8
4s=2s+22
2s=22
s=11
b=18
m=36
f=44
The wife's age is 36 years old.